OpenStudy (anonymous):

Based on the graph of y = x2 − 2x − 3, what is the x coordinate of the positive x-intercept?

OpenStudy (anonymous):

Solve this equation and you will get two answers. The one with + sign is the right one :)

OpenStudy (anonymous):

Do you know how to solve it?

OpenStudy (anonymous):

You could factor it or graph it, but it wants you to graph it

OpenStudy (anonymous):

i still don't know how to graph it

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

Do u know the shape of x^2 graph?

OpenStudy (anonymous):

y = ax2 + bx + c since a i positive, this parabola will be opening upward First we want to find the vertex of this parabola. We can find the x coordinate by the formula -b/2a, in our equation b= -2 and a=1 so we have 2/2*1 which means the x coordinate is 1, to find the y coordinate we plug in 1 to the equation, y=1^2-2(1)-3 Y=1-2-3=-4 the coordinates of the vertex is (1,-4) now you can make an x and y table to find the other points.